Grade 11General Physics

If area velocity and density are fundamental units then dimensional formula of force is

dm kathi
8 Years agoGrade 11

1 Answer

Aashu Mittal
8 Years ago
Let write Area as `A`, Velocity as`V` and density as `D`Now, We know that, A=[L^2], V=[LT^-1] and D=[ML^-3]We also know that dimensional formula of force is F=[MLT^-2]So , we only have do is to obtain dimensional formula of force from dimensional formulae of A, V and D.Now if we multiply A and D , we will get [AD]=[L^2][ML^-3]=[ML^-1]Now squaring V and multiplying it with AD, we will get, [V^2][AD]=[L^2T^-2][ML^-1]=[MLT^-2]This is dimensional formula of force so answer is [AV^2D].......Answer.